# A Minimalist Optimizer Design for LLM Pretraining

### v2

@@ A.9 Discussion on why Column-wise matters for the Last Layer
and (A.2), respectively.
Now plugging this back to (A.9) we get:
$\mathbb{E}\left[\phi^{t+1}\right]-\mathbb{E}\left[\phi^{t}\right]$ (A.11)
$\leq$ $\rho_{0}s\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\right)\sigma_{l}^{2}+\sum_{l=1}^{L}\frac{\gamma\eta_{l}^{2}}{2}\sigma_{l}^{2}+\sum_{l=1}^{L}2c_{l,1}\eta_{l}^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}$
$+\sum_{l=1}^{L}(-\eta_{l}+\frac{\eta_{l}}{2\rho_{0}}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{l,1}\eta_{l}^{2})\mathbb{E}\|\nabla_{l}^{t}\|^{2}+2\rho_{0}\left(\sum_{l=1}^{L}\eta_{l}\right)\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\mathbb{E}\left[\left\|\nabla_{l}^{t}\right\|^{2}\right]\right)$
$+4\sum_{i=1}^{t-1}\sum_{l=1}^{L}c_{l,1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}a_{l,t,i}\mathbb{E}\left[\left\|\theta^{i+1}-\theta^{i}\right\|^{2}\right]$
@@
$+\sum_{i=1}^{t-1}\sum_{l=1}^{L}(c_{l,i+1}-c_{l,i})\mathbb{E}\left[\left\|\theta_{l}^{t+1-i}-\theta_{l}^{t-i}\right\|^{2}\right]$
@@
$c_{l,i+1}\leq c_{l,i}-\left(4c_{l,1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1-\beta_{l}^{t}}{1-\beta_{l}})^{2}\right)a_{l,t,t-i}$
for all l=1,,Ll=1,...,L.
@@
$c_{l,i+1}=c_{l,i}-\left(4c_{l,1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\gamma^{2}\beta_{l}^{i}\left(i+\frac{\beta_{l}}{1-\beta_{l}}\right)$
also we can take $c_{l,1}$ such that
$c_{l,1}=$ $\sum_{i=1}^{\infty}\left(4c_{l,1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\gamma^{2}\beta_{l}^{i}\left(i+\frac{\beta_{l}}{1-\beta_{l}}\right)$
$=$ $\left(4c_{l,1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\gamma^{2}\left(\sum_{i=1}^{\infty}i\beta_{l}^{i}+\frac{\beta_{l}}{1-\beta_{l}}\sum_{i=1}^{\infty}\beta_{l}^{i}\right)$
$=$ $\left(4c_{l,1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\frac{\gamma^{2}\beta_{l}(1+\beta_{l})}{(1-\beta_{l})^{2}}$
@@
$c_{l,1}=\frac{2\rho_{0}s\eta_{l}^{2}\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}}{1-4\eta_{l}^{2}\frac{\gamma^{2}\beta_{l}(1+\beta_{l})}{(1-\beta_{l})^{2}}}.$ (A.12)
Note that we can give an upper bound for $c_{l,1}$ by requiring the denominator $\geq 1/2$, i.e.
$\eta_{l}\leq\frac{1-\beta_{l}}{\gamma\sqrt{8\beta_{l}(1+\beta_{l})}},$ (A.13)
@@
$c_{l,1}\leq 4\rho_{0}s\eta_{l}^{2}\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}.$ (A.14)
Now we have (since the last three terms of (A.11) sum to negative)
$\mathbb{E}\left[\phi^{t+1}\right]-\mathbb{E}\left[\phi^{t}\right]$ (A.15)
$\leq$ $\rho_{0}s\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\right)\sigma_{l}^{2}+\sum_{l=1}^{L}\frac{\gamma\eta_{l}^{2}}{2}\sigma_{l}^{2}+\sum_{l=1}^{L}2c_{l,1}\eta_{l}^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}$
$+\sum_{l=1}^{L}\left(-\eta_{l}+\frac{\eta_{l}}{2\rho_{0}}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{l,1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\right)\mathbb{E}\|\nabla_{l}^{t}\|^{2}$
@@
$\frac{3}{4}\eta_{l}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{l,1}\eta_{l}^{2}+4s\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\leq-\frac{\eta_{l}}{2},$ (A.16)
the coefficient of $\mathbb{E}\|\nabla_{l}^{t}\|$ can be greatly simplified. Note that we can guarantee (A.16) if
$\eta_{l}\leq\min\left\{\frac{1}{8\gamma},\frac{1}{8\gamma L}(\frac{1-\beta_{l}}{\beta_{l}})^{2},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{L\beta_{l}(1+\beta_{l})}}\right\}$ (A.17)
for all l=1,2,,Ll=1,2,...,L. Here each term in the right hand side of (A.17) is bounding each term on the left hand side of (A.16).
@@
$\sum_{l=1}^{L}\frac{\eta_{l}}{2}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq\mathbb{E}\left[\phi^{t}\right]-\mathbb{E}\left[\phi^{t+1}\right]$ (A.18)
$+2s\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\right)\sigma_{l}^{2}+\sum_{l=1}^{L}\frac{\gamma\eta_{l}^{2}}{2}\sigma_{l}^{2}+\sum_{l=1}^{L}2c_{l,1}\eta_{l}^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}.$
Note that we have defined $s:=\sum_{l=1}^{L}\eta_{l}$. We can upper bound it by $s\leq L/\sqrt{\gamma T}$ if we assume $\eta_{l}\leq 1/\sqrt{\gamma T}$. Combining (A.13) and (A.17) we get:
$\eta_{l}\leq\min\left\{\frac{1}{8\gamma},\frac{1}{8\gamma L}(\frac{1-\beta_{l}}{\beta_{l}})^{2},\frac{1-\beta_{l}}{\gamma\sqrt{8\beta_{l}(1+\beta_{l})}},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{L\beta_{l}(1+\beta_{l})}},\frac{1}{\sqrt{\gamma T}}\right\},$ (A.19)
@@
$\eta_{l}=\min\left\{\frac{1}{8\gamma},\frac{1}{8\gamma L}(\frac{1-\beta_{l}}{\beta_{l}})^{2},\frac{1-\beta_{l}}{4\gamma},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{2L}},\frac{1}{\sqrt{\gamma T}}\right\}.$ (A.20)
@@
$\eta_{l}^{2}\leq\frac{1}{8\gamma L}(\frac{1-\beta_{l}}{\beta_{l}})^{2}\frac{1}{\sqrt{\gamma T}}.$ (A.21)
Plugging (A.21) and (A.14) into (A.18), we obtain:
$\sum_{l=1}^{L}\eta_{l}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq$ $2(\mathbb{E}\left[\phi^{t}\right]-\mathbb{E}\left[\phi^{t+1}\right])$ (A.22)
@@
$\eta_{l}$ $=\min\left\{\frac{1}{8\gamma},\frac{1}{8\gamma L}(\frac{1-\beta_{l}}{\beta_{l}})^{2},\frac{1-\beta_{l}}{4\gamma},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{2L}}\right\}$
$\geq\min\left\{\frac{1}{8\gamma},\frac{\delta^{2}}{8\gamma L},\frac{\delta}{4\gamma},\frac{\delta}{4\gamma}\sqrt[3]{\frac{\delta}{2L}},\frac{1}{\sqrt{\gamma T}}\right\}=:\eta$ (A.23)
@@
$\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq$ $\frac{2(\mathbb{E}\left[\phi^{t}\right]-\mathbb{E}\left[\phi^{t+1}\right])}{\eta}$ (A.24)
@@
Now by telescoping sum of (A.22) for t=1,,Tt=1,...,T, also notice that $1/\eta=\Theta(\sqrt{\gamma T}L\gamma/\delta^{2})$ as $T$ grows, we get our final result of
$\frac{1}{T}\sum_{t=1}^{T}\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq\frac{2L\gamma^{3/2}\mathbb{E}\Delta_{1}}{\delta^{2}\sqrt{T}}+\sum_{l=1}^{L}\left(\frac{1-\beta_{l}}{1+\beta_{l}}\frac{L\sqrt{\gamma}}{4\sqrt{T}}+\frac{L\gamma^{3/2}}{2\sqrt{T}}+\frac{1-\beta_{l}}{\beta_{l}^{3}}\frac{\gamma^{2}}{4LT}\right)\frac{\sigma_{l}^{2}}{\delta^{2}}.$ (A.25)

### v3

@@ Appendix I Fine-tuning Results
and (13), respectively.
Now plugging this back to (20) we get:
$\mathbb{E}\left[\phi^{t+1}\right]-\mathbb{E}\left[\phi^{t}\right]$ (22)
$\leq$ $\rho_{0}s\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\right)\sigma_{l}^{2}+\sum_{l=1}^{L}\frac{\gamma\eta_{l}^{2}}{2}\sigma_{l}^{2}+\sum_{l=1}^{L}2c_{1}\eta_{l}^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}$
$+\sum_{l=1}^{L}(-\eta_{l}+\frac{\eta_{l}}{2\rho_{0}}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{1}\eta_{l}^{2})\mathbb{E}\|\nabla_{l}^{t}\|^{2}+2\rho_{0}\left(\sum_{l=1}^{L}\eta_{l}\right)\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\mathbb{E}\left[\left\|\nabla_{l}^{t}\right\|^{2}\right]\right)$
$+4\sum_{i=1}^{t-1}\sum_{l=1}^{L}c_{1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}a_{l,t,i}\mathbb{E}\left[\left\|\theta^{i+1}-\theta^{i}\right\|^{2}\right]$
@@
$+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}\left[\left\|\theta^{t+1-i}-\theta^{t-i}\right\|^{2}\right]$
@@
$c_{i+1}\leq c_{i}-\sum_{l=1}^{L}\left(4c_{1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1-\beta_{l}^{t}}{1-\beta_{l}})^{2}\right)a_{l,t,t-i}.$
@@
$c_{i+1}=c_{i}-\sum_{l=1}^{L}\left(4c_{1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\gamma^{2}\beta_{l}^{i}\left(i+\frac{\beta_{l}}{1-\beta_{l}}\right)$
also we can take $c_{1}$ such that
$c_{1}=$ $\sum_{l=1}^{L}\sum_{i=1}^{\infty}\left(4c_{1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\gamma^{2}\beta_{l}^{i}\left(i+\frac{\beta_{l}}{1-\beta_{l}}\right)$
$=$ $\sum_{l=1}^{L}\left(4c_{1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\gamma^{2}\left(\sum_{i=1}^{\infty}i\beta_{l}^{i}+\frac{\beta_{l}}{1-\beta_{l}}\sum_{i=1}^{\infty}\beta_{l}^{i}\right)$
$=$ $\sum_{l=1}^{L}\left(4c_{1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}\gamma^{2}(\frac{1}{1-\beta_{l}})^{2}\right)\frac{\gamma^{2}\beta_{l}(1+\beta_{l})}{(1-\beta_{l})^{2}}$
@@
$c_{1}=\frac{\sum_{l=1}^{L}2\rho_{0}s\eta_{l}^{2}\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}}{1-4\sum_{l=1}^{L}\eta_{l}^{2}\frac{\gamma^{2}\beta_{l}(1+\beta_{l})}{(1-\beta_{l})^{2}}}.$ (23)
Note that we can give an upper bound for $c_{1}$ by requiring the denominator $\geq 1/2$, i.e.
$\eta_{l}\leq\frac{1-\beta_{l}}{\gamma\sqrt{8\beta_{l}(1+\beta_{l})}L},\ \forall l=1,2,...,L,$ (24)
@@
$\sum_{l=1}^{L}2\rho_{0}s\eta_{l}^{2}\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}\leq c_{1}\leq\sum_{l=1}^{L}4\rho_{0}s\eta_{l}^{2}\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}.$ (25)
Now we have (since the last three terms of (22) sum to negative)
$\mathbb{E}\left[\phi^{t+1}\right]-\mathbb{E}\left[\phi^{t}\right]$ (26)
$\leq$ $\rho_{0}s\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\right)\sigma_{l}^{2}+\sum_{l=1}^{L}\frac{\gamma\eta_{l}^{2}}{2}\sigma_{l}^{2}+\sum_{l=1}^{L}2c_{1}\eta_{l}^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}$
$+\sum_{l=1}^{L}\left(-\eta_{l}+\frac{\eta_{l}}{2\rho_{0}}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{1}\eta_{l}^{2}+2\rho_{0}s\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\right)\mathbb{E}\|\nabla_{l}^{t}\|^{2}$
@@
$\frac{3}{4}\eta_{l}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{1}\eta_{l}^{2}+4s\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\leq-\frac{\eta_{l}}{2},$ (27)
the coefficient of $\mathbb{E}\|\nabla_{l}^{t}\|$ can be greatly simplified. Note that we can guarantee (27) if
$\eta_{l}\leq\min\left\{\frac{1}{8\gamma},\frac{1}{64\sum_{l=1}^{L}\rho_{0}s\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{L\beta_{l}(1+\beta_{l})}}\right\}$ (28)
for all l=1,2,,Ll=1,2,...,L. Here each term in the right hand side of (28) is bounding each term on the left hand side of (27).
@@
$\sum_{l=1}^{L}\frac{\eta_{l}}{2}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq\mathbb{E}\left[\phi^{t}\right]-\mathbb{E}\left[\phi^{t+1}\right]$ (29)
$+2s\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\right)\sigma_{l}^{2}+\sum_{l=1}^{L}\frac{\gamma\eta_{l}^{2}}{2}\sigma_{l}^{2}+\sum_{l=1}^{L}2c_{1}\eta_{l}^{2}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}.$
Note that we have defined $s:=\sum_{l=1}^{L}\eta_{l}$. We can upper bound it by $s\leq L/\sqrt{\gamma T}$ if we assume $\eta_{l}\leq 1/\sqrt{\gamma T}$. Combining (24) and (28) we get:
$\eta_{l}\leq\min\left\{\frac{1}{8\gamma},\frac{1}{8\gamma L}(\frac{1-\beta_{l}}{\beta_{l}})^{2},\frac{1-\beta_{l}}{\gamma\sqrt{8\beta_{l}(1+\beta_{l})}},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{L\beta_{l}(1+\beta_{l})}},\frac{1}{\sqrt{\gamma T}}\right\},$ (30)
@@
$\eta_{l}=\min\left\{\frac{1}{8\gamma},\frac{1}{64\sum_{l=1}^{L}\rho_{0}s\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}},\frac{1-\beta_{l}}{4\gamma},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{2L}},\frac{1}{\sqrt{\gamma T}}\right\}.$ (31)
@@
$\eta_{l}^{2}\leq\frac{1}{8\gamma L}(\frac{1-\beta_{l}}{\beta_{l}})^{2}\frac{1}{\sqrt{\gamma T}}.$ (32)
Plugging (32) and (25) into (29), we obtain:
$\sum_{l=1}^{L}\eta_{l}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq$ $2(\mathbb{E}\left[\phi^{t}\right]-\mathbb{E}\left[\phi^{t+1}\right])$ (33)
@@
$\eta_{l}$ $=\min\left\{\frac{1}{8\gamma},\frac{1}{64\sum_{l=1}^{L}\rho_{0}s\gamma^{4}\beta_{l}\frac{1+\beta_{l}}{(1-\beta_{l})^{4}}},\frac{1-\beta_{l}}{4\gamma},\frac{1-\beta_{l}}{4\gamma}\sqrt[3]{\frac{1-\beta_{l}}{2L}}\right\}$
$\geq\min\left\{\frac{1}{8\gamma},\frac{\delta^{4}}{64L\rho_{0}s\gamma^{4}},\frac{\delta}{4\gamma},\frac{\delta}{4\gamma}\sqrt[3]{\frac{\delta}{2L}},\frac{1}{\sqrt{\gamma T}}\right\}=:\eta$ (34)
@@
$\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq$ $\frac{2(\mathbb{E}\left[\phi^{t}\right]-\mathbb{E}\left[\phi^{t+1}\right])}{\eta}$ (35)
@@
Now by telescoping sum of (33) for t=1,,Tt=1,...,T, also notice that $1/\eta=\Theta(\sqrt{\gamma T}L\gamma/\delta^{2})$ as $T$ grows, we get our final result of
$\frac{1}{T}\sum_{t=1}^{T}\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq\frac{2L\gamma^{3/2}\mathbb{E}\Delta_{1}}{\delta^{2}\sqrt{T}}+\sum_{l=1}^{L}\left(\frac{1-\beta_{l}}{1+\beta_{l}}\frac{L\sqrt{\gamma}}{4\sqrt{T}}+\frac{L\gamma^{3/2}}{2\sqrt{T}}+\frac{1-\beta_{l}}{\beta_{l}^{3}}\frac{\gamma^{2}}{4LT}\right)\frac{\sigma_{l}^{2}}{\delta^{2}}.$ (36)